OpenStudy (zolock):

If f(x) = 9 − 4|x + 1|, find f(−5). (Points : 1) a. −15 b. −7 c. 7 d. 23

OpenStudy (johnweldon1993):

See how the only difference is f(x) is now f(-5) ... the 'x' was replaced by -5 ...we will do that in the equation too! Replace every 'x' you see with -5 f(-5) = 9 - 4|-5 + 1| Deal with the absolute value first....-5+ 1 = -4 right? well the absolute value of that is 4...so f(-5) = 9 - 4(4) Can you take it from here?

OpenStudy (zolock):

whats the answer

OpenStudy (johnweldon1993):

What's 9 - 4(4) ?

OpenStudy (zolock):

-7

OpenStudy (johnweldon1993):

There you go!

OpenStudy (zolock):

thank u

OpenStudy (zolock):

can u help with more ?s

OpenStudy (zolock):

If f(x) = |x + 6|, find f(−10). (Points : 1) a. −16 b. 16 c. −4 d. 4